OpenStudy (anonymous):

How do I simplify ((z^p)^7)/(z^7)^(p-1)?

OpenStudy (anonymous):

help meeeeeeeee

OpenStudy (zzr0ck3r):

\[\frac{(z^p)^7}{(z^7)^{p-1}}\]?

OpenStudy (zzr0ck3r):

\[\frac{z^{7p}}{z^{7p-7}}=z^{7p-(7p-7)}=z^{7p-7p+7}=z^7\]

OpenStudy (cggurumanjunath):

z^7

OpenStudy (ybarrap):

Note, \( z^{a+b} = z^az^b\) and \(1/z^{a} = z^{-a}\) and \(z^{an}=(z^{a} )^{n} \). Using just these 3 rules you can solve this mildly intimidating problem.